OpenStudy (anonymous):

Hey people! Obama won :D! No but seriously could someone help me with this question? If you were a Romney supporter and you know this question..... Can you please still help me? Jk jk no hard feeling :) So the question is: Create your own unique quadratic equation in the form y = ax^2 + bx + c that opens the same direction and shares one of the x-intercepts of the graph of y = x^2 - 10x + 16. Determine the following. Explain whether the graph has a maximum or minimum. Find the vertex and x-intercepts of the graph.

OpenStudy (anonymous):

Come on! I had the election PVRed. Now you've gone and spoiled the ending.

OpenStudy (anonymous):

Oh my gosh! I'm am so sorry...... So who were you rooting for? Obama or Romney?

OpenStudy (anonymous):

MERICA!!!!

OpenStudy (anonymous):

I'm just kidding. I'm Canadian, so I supported whichever one was in support of boosting igloo development, eh?

OpenStudy (anonymous):

hahaha ok good :)...... so do you by chance know the answer to my math question?

OpenStudy (anonymous):

going up since a is bigger than o

OpenStudy (anonymous):

since it's going up so it got a local min

OpenStudy (anonymous):

x-intercepts is the root

OpenStudy (anonymous):

means when y=0

OpenStudy (anonymous):

I'm sorry but I'm confused......

OpenStudy (anonymous):

when the local mins should be -b/2a b=-10 a=1

OpenStudy (anonymous):

or you could think that the local min or local max is exactly between the two root

OpenStudy (anonymous):

I'm sorry but I dot understand, what your saying is really confusing...... Could you answer the whole question instead of bits and pieces? Its just hard to understand what your trying to say......

OpenStudy (lgbasallote):

i'm a clinton supporter

OpenStudy (anonymous):

congrats but I needed help.........

OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME! I AM TRYING TO ACTUALLY GO TO BED TONIGHT!!!

OpenStudy (anonymous):

ok, for the equation x^2 - 10x + 16 a=1 b=-10 c=16, and a>0, so it's concave up, and there only one chance for a local min. what's the value of local min? the equation for local min or local max is -b/2a=10/2=5, and when x=5, y=25-50+16=-9 so the local min point is (5,-9)

OpenStudy (anonymous):

which part do you not understand know?